OpenStudy (anonymous):

Find the antiderivative for (dP/dx)=36-6x

OpenStudy (anonymous):

\[\frac{dP}{dx}=36-6x\implies P=\int\limits\limits(36-6x)dx=36x-3x^2\]

OpenStudy (anonymous):

P = 36x - 3x^2 + constant

OpenStudy (anonymous):

Yep, forgot the constant :P

OpenStudy (anonymous):

thanks!! :D